OpenStudy (anonymous):

Find the point on the graph of z=−x^2+y^2 at which vector n=<4,−6,1> is normal to the tangent plane.

OpenStudy (anonymous):

How do I approach this? I'm completely lost right now!

OpenStudy (amistre64):

hmm i beleive the derivative of a surface IS the normal vector equations

OpenStudy (amistre64):

\[D[z=-x^2+y^2]\to <\frac{dz}{dx},\frac{dz}{dy},\frac{dz}{dz}>\] \[<-2x, 2y, 1>\]

OpenStudy (amistre64):

so when x=-2 and y = 3, we should have it

OpenStudy (anonymous):

Nope. WeBWorK tells me that is not the correct answer.

OpenStudy (amistre64):

then ill have to review how to find a gradient of a surface ...

OpenStudy (amistre64):

http://tutorial.math.lamar.edu/Classes/CalcIII/GradientVectorTangentPlane.aspx last problem i believe is the same structure is this one

OpenStudy (amistre64):

y = -3 might be better since we are trying to get a -6 from 2y

OpenStudy (amistre64):

plug in for x=-2, and y=-3 to find the value for z then your point is: (-2,-3,z)

OpenStudy (anonymous):

I tried that with z=1, and it told me that -3 is correct for the y-coordinate, but the other two were still wrong. Changing the sign on the x-coordinate gave me two correct coordinates, but I still don't know how to get the z-coordinate.

OpenStudy (amistre64):

you are given that: z = -x^2 + y^2 if you know x and y, plug them in to get "z" :)

OpenStudy (anonymous):

Oh...duh!

OpenStudy (amistre64):

-2x = 4; when x=-2 so either your problem is entered wrong on here or ther eis a misprint in your material

OpenStudy (anonymous):

I copied it directly, so I'm guessing it's wrong on the server side of WeBWorK. Thank you so much though!

OpenStudy (amistre64):

youre welcome :) and good luck